Identify each of the equations as representing a circle, a parabola, an ellipse, a hyperbola, or none of these:
a) x(x-3)=y(1-2y^2)
b)2x^2+4y^2=y+2x
c)y=3(1-2x)(1+2x)

Answer by stanbon(75887)

Identify each of the equations as representing a circle, a parabola, an ellipse, a hyperbola, or none of these:
a) x(x-3)=y(1-2y^2)
none of these because it is a cubic
-----------------------------------------
b)2x^2+4y^2=y+2x
ellipse because you have x^2 and y^2 and their coefficients are positive
and not equal.
--------------------------------
c)y=3(1-2x)(1+2x)
parabola because it is a quadratic with independent variable "x".
===================================================
Cheers,
Stan H.